### v3

@@ 3.4 Inference for associations between two vector time series
Formal inference methods for $\Omega$ based on its LaDynS estimate (Eq. 10) are not available, but because LaDynS reduces to graphical LASSO (gLASSO) when the weights $w_{k}^{(t)}$ in Eq. 11 are fixed, we co-opt gLASSO inference methods. Specifically, Jankova et al. (2015) suggested de-sparsifying the gLASSO estimate $\widehat{\Omega}$ according to
@@
We applied this result to the de-sparsified LaDynS estimate of $\Omega$, even though we do not quite have a gLASSO setup, and we verified by simulation that its elements are indeed approximately normal in Section 4.2.2. Jankova et al. (2015) also proposed an estimator of $\operatorname{Var}[\widetilde{\Omega}_{12}^{(t,s)}]$ that is likely to be downward biased in our framework since estimating the canonical weights $w_{k}^{(t)}$ induces extra randomness. Instead, we use the bootstrap estimate $\widehat{\operatorname{Var}}[\widetilde{\Omega}_{12}^{(t,s)}]$ described at the end of this section. We then rely on Eq. 14 to obtain p-values
@@
below a pre-specified level $\alpha_{\text{BH}}$ using the procedure of Benjamini and Hochberg (1995) (BH). To proceed, let ${p}^{[1]}\leq\dots\leq{p}^{[n_{\text{roi}}]}$ denote the ordered permutation bootstrap p-values ${p}^{(t,s)}$ that correspond to $n_{\text{roi}}$ cross-precision elements in the region of interest. Then, we find the maximum $k_{\text{BH}}$ satisfying ${p}^{[k_{\text{BH}}]}\leq\frac{k_{\text{BH}}}{n_{\text{roi}}}\alpha_{\text{BH}}$ and reject $H_{0}^{(t,s)}$ with ${p}^{(t,s)}$ smaller than $\frac{k_{\text{BH}}}{n_{\text{roi}}}\alpha_{\text{BH}}$. The FDR guarantee is established by Benjamini and Hochberg (1995) as long as the ${p}^{(t,s)}$’s are independent and valid p-values.
Cluster-wise inference by excursion test: As a further safeguard against falsely detecting correlated activity between brain areas, we obtain p-values for each identified connectivity epoch using the excursion test of Ventura et al. (2005), as follows. For each cluster $k$ identified by the BH procedure, we calculate the test statistic:
@@
which is reminiscent of Fisher’s method for testing the global significance of multiple hypotheses (Fisher, 1925). Large values of $T_{k}$ provide evidence against cross-area connectivity in cluster $k$, so we calculate the corresponding p-value as $\int_{T_{k}}^{\infty}f_{T_{\max}}(t)~dt$, where $f_{T_{\max}}$ is the null distribution of $T_{\max}:=\max_{j}T_{j}$ under the global null hypothesis of no connectivity anywhere. We use $f_{T_{\max}}$ rather than the respective null distributions of each $T_{k}$ to control the family-wise type I error rather than the type I error for each cluster. We approximate $f_{T_{\max}}$ by the previous permutation bootstrap: for each permuted dataset $b=1,\ldots,B$, we estimate the cross-precision matrix and corresponding p-values, identify all clusters of p-values below $\frac{k_{\text{BH}}}{n_{\text{roi}}}\alpha_{\text{BH}}$, calculate the corresponding test statistics in Eq. 17, and let $T_{\max}^{b}$ be their maximum. The $B$ values $T_{\max}^{b}$ are samples from $f_{T_{\max}}$, which we use to approximate the p-value for cluster $k$ by the sampling proportion:
@@ 3.5 Locally Stationary State-space Model and Local Granger Causality
This state-space formulation is convenient to impose local stationarity on the latent time series, which we do by fitting stationary state-space models in moving windows of time. Local stationarity is justified because the functional connectivity within and between brain regions changes relatively slowly over time. The state-space formulation is also convenient to calculate the Granger causality between regions: $Z_{2}$ is said to Granger-cause $Z_{1}$ at time $t$ if some $\alpha_{12,s}^{(t)}$ are non-zero (Ombao and Pinto, 2022) (and conversely if some $\alpha_{21,s}^{(t)}$ are non-zero). The coefficient of partial determination (partial $R^{2}$) between $\left(Z_{2}^{(t-d_{\mathrm{cross}})},\dots,Z_{2}^{(t-1)}\right)$ and $Z_{1}^{(t)}$, conditional on $Z_{1}^{(t-d_{\mathrm{auto}})},\dots,Z_{1}^{(t-1)}$, calculated as
@@
where $\tau_{2}$ is the maximum connection time from brain regions 2 to 1, $\tau_{1}\leq\tau_{2}\leq d_{\mathrm{cross}}$. We set $\tau_{2}=d_{\mathrm{cross}}$ by default, unless there is reason to consider shorter connection times. A plug-in estimator of $R^{2}_{2\rightarrow 1}(t)$ is easily obtained from the estimated covariance matrix of $\left(Z_{1}^{(t)},\dots,Z_{1}^{(t-d_{\mathrm{auto}})},Z_{2}^{(t-1)},\dots,Z_{2}^{(t-d_{\mathrm{cross}})}\right),$ without actually running the regressions (Anderson-Sprecher, 1994).
@@ 4 Results
We have introduced LaDynS to estimate the dynamic connectivity between two or more multivariate time series, and we proposed inference procedures to identify when connectivity is statistically significant. We apply LaDynS to experimental data in Section 4.4, but first we present the theoretical results on the equivalence between generative pCCA models and model-free CCA methods. In particular, Theorem 4.3 establishes an equivalence between the GENVAR version of multi-set CCA (Kettenring, 1971) and maximum likelihood applied to our dynamic pCCA. Next we examine its performance on simulated data that have properties similar to the experimental data. The data are simulated from the shared oscillatory driver model described in Section 4.2, which is unrelated to the LaDynS model. Section S3.2 contains results based on simulated data that are consistent with the LaDynS model. In Section 4.3, we compare the performance of LaDynS to other existing methods. The reproducible code scripts for the simulations and experimental data analyses are provided at github.com/HeejongBong/ladyns.
@@ 4.1 Equivalence between dynamic pCCA model and multiset CCA
We first revisit the connection between classical CCA and the probabilistic CCA formulation of Bach and Jordan (2005). Let $\{(X_{1[n]},X_{2[n]})\}_{n=1}^{N}$ be independent observations drawn from the joint distribution in Eq. 2, and denote the maximum likelihood estimates by $(\widehat{\beta}_{1},\widehat{\beta}_{2})$. The result below restates Theorem 2 of Bach and Jordan (2005).
Theorem 4.1 (Bach and Jordan, 2005, Theorem 2).
@@
We now derive a similar connection between the multiset generalization of CCA introduced by Kettenring (1971) and the dynamic pCCA model in Eqs. 5 and 6. The proof is provided in Appendix S1.
@@ 4.2 LaDynS performance on simulated data from a shared oscillatory driver model with time delay
We used a probabilistic model of the inter-areal coherence in local field potentials to generate LFP time series with a dynamic lead-lag relationship between two brain areas. The shared oscillatory driver model with time delay in Ombao and Pinto (2022) assumes that the coherence between two univariate LFP time series $L_{1}$ and $L_{2}$ is driven by a univariate latent oscillation $L_{0}$ at frequency $f_{0}$:
@@
where $\eta_{k}^{(t)}$ is regional baseline noise, and $\tau_{k}$ is the lead-lag from $L_{0}$ to $L_{k}$. This model generalizes the Synaptic-Source-Mixing (SSM) model of Schneider et al. (2021), which was shown to be capable of modeling the dynamic cross-regional coherence observed in experimental LFP data between frontal and parietal cortices, as well as between LGN and the visual cortex.
For our simulation, we extended Eq. 26 to generate two sets of multi-dimensional LFP time series $L_{k}$ of dimension $d_{k}$, $k=1,2$, and we used three latent oscillations $L_{0,j}$, $j=1,2,3$, to allow non-stationary lead-lag relationships between the two brain areas. The resulting model was a special case of the Latent Dynamic Factor Analysis model in Bong et al. (2020):
@@
where the factor loadings $\beta_{kj}\in\mathbb{R}^{d_{k}}$ were vectors of dimension $d_{k}=25$. We virtually arranged the $25$ components of $L_{1}^{(t)}$ and $L_{2}^{(t)}$ on $5\times 5$ regular two-dimensional arrays so we could create simulated data exhibiting spatial correlations similar to those in the experimental data in Section 4.4. To do that we let $\eta_{k}^{(t)}$ be pink noise, as in Schneider et al. (2021), which is a mixture of oscillations at different frequencies $f$ with power proportional to $f^{-\alpha}$ – we used $\alpha=1.4$ – and we added the spatial correlations $\operatorname{Cov}[\eta_{ki}^{(t)},\eta_{kj}^{(t)}]=\exp(-\frac{\mathrm{dist}^{2}(i,j)}{2\sigma_{\mathrm{spatial}}^{2}})$, where $\mathrm{dist}(i,j)$ was the distance between electrodes $i$ and $j$ on one array and $\sigma_{\text{spatial}}=0.8$. Next, we let the factor loadings $\beta_{kj}$ have components $\beta_{kji}=\gamma\cdot\exp(-\frac{\mathrm{dist}^{2}(i,p_{kj})}{2\sigma_{\mathrm{spatial}}^{2}})$, $i=1,\ldots d_{k}$, where $p_{kj}$ was a randomly chosen location on an array and $\gamma$ took values in $\{0.055,0.063,0.071,0.077,0.084,0.089\}$. This set spanned $\{0.5,0.66,0.84,1.0,1.16,1.34\}$ times the observed signal-to-noise ratio (SNR) in the experimental data in Section 4.4. Finally, the three latent oscillations $L_{0,j}$ were set at $18$ Hz and designed to induce three epochs of lead-lag relationships between $L_{1}$ and $L_{2}$ around experimental times $80$, $200$ and $400$ ms; $L_{1}$ lead $L_{2}$ by $30$ ms during the first epoch, and $L_{1}$ lagged $L_{2}$ by $30$ ms during the other two, as depicted in Figure 5(a).
One simulated dataset consisted of $N=1000$ trials of $500$ ms long time series $\{L_{1}^{(t)}\in\mathbb{R}^{d_{1}},L_{2}^{(t)}\in\mathbb{R}^{d_{2}}\}$ generated from Eq. 27 at sampling frequency $1000$ Hz. We filtered the simulated LFP recordings using the complex Morlet wavelet at frequency $f_{0}=18$ Hz and bandwidth $50$ ms, the same we applied to the data in Section 4.4, and collected the beta oscillation amplitude envelopes as the absolute values of the filtered signals. The complex Morlet wavelet is a complex sinusoidal with a Gaussian envelope (Gabor, 1946), where the bandwidth refers to the Gaussian standard deviation. (The filtered amplitude at a given time $t$ is essentially the amplitude that would be obtained using a harmonic regression with cosine and sine damped by the Gaussian kernel centered at $t$; the wavelet formulation is computationally different and more efficient.) After downsampling the power envelopes to $100$ Hz, we applied LaDynS to the resulting data $X_{1}$ and $X_{2}$ with $T=50$ time points.
@@ 4.2.2 Results
Figures 5 and 6 used connection strength $\gamma=0.077$. We repeated the simulation for a range of $\gamma$ such that the SNR in the simulated data beta band ranged from zero to twice the estimated SNR of the experimental data in Section 4.4. Instead of showing graphs we summarized the performance of LaDynS using false cluster discovery and non-discovery rates (Perone Pacifico et al., 2004) (FCDR, FCNR) in the estimated precision matrix. We defined a cluster of estimated nonzero precision elements to be falsely discovered if it contained no true effect, and a true cluster was deemed falsely non-discovered if no estimated cluster overlapped with it. We estimated these rates with the corresponding proportions in $60$ simulated datasets. They are plotted in Figure 7 against nominal significance levels of the excursion test between 0 and 10%, where the target FDR for the BH procedure was fixed at 5%. The FCDR remained below the nominal level across all tested significance values and connectivity intensities (all or almost all discovered clusters were true clusters) and for large enough $\gamma$, the FCNR was zero (all clusters were discovered).

### v4

@@ 3.4 Inference for associations between two vector time series
Formal inference methods for $\Omega$ based on its LaDynS estimate (Eq. 10) are not available, but because LaDynS reduces to graphical LASSO (gLASSO) when the weights $w_{k}^{(t)}$ in Eq. 11 are fixed, we co-opt gLASSO inference methods. Specifically, jankova2015confidence suggested de-sparsifying the gLASSO estimate $\widehat{\Omega}$ according to
@@
The assumptions include independence across experimental trials, irrepresentability condition and bounded eigenvalues of $\Omega$ (see Assumptions (A1) and (A2) in jankova2015confidence). While some of these conditions may be strong and difficult to verify in practice, they are commonly imposed in the literature on sparse graph estimation with $\ell_{1}$ regularization. In Section S3.2, we examine the asymptotic normality of the de-sparsified LaDynS estimate using simulation. We applied this result to the de-sparsified LaDynS estimate of $\Omega$, even though we do not quite have a gLASSO setup, and we verified by simulation that its elements are indeed approximately normal in Section 4.2.2. jankova2015confidence also proposed an estimator of $\operatorname{Var}[\widetilde{\Omega}_{12}^{(t,s)}]$ that is likely to be downward biased in our framework since estimating the canonical weights $w_{k}^{(t)}$ induces extra randomness. Instead, we use the bootstrap estimate $\widehat{\operatorname{Var}}[\widetilde{\Omega}_{12}^{(t,s)}]$ described at the end of this section. We then rely on Eq. 14 to obtain p-values
@@
below a pre-specified level $\alpha_{\text{BH}}$ using the procedure of benjamini1995controlling (BH). To proceed, let ${p}^{[1]}\leq\dots\leq{p}^{[n_{\text{roi}}]}$ denote the ordered permutation bootstrap p-values ${p}^{(t,s)}$ that correspond to $n_{\text{roi}}$ cross-precision elements in the region of interest. Then, we find the maximum $k_{\text{BH}}$ satisfying ${p}^{[k_{\text{BH}}]}\leq\frac{k_{\text{BH}}}{n_{\text{roi}}}\alpha_{\text{BH}}$ and reject $H_{0}^{(t,s)}$ with ${p}^{(t,s)}$ smaller than $\frac{k_{\text{BH}}}{n_{\text{roi}}}\alpha_{\text{BH}}$. The FDR guarantee is established by benjamini1995controlling as long as the ${p}^{(t,s)}$’s are independent and valid p-values.
Cluster-wise inference by excursion test: As a further safeguard against falsely detecting correlated activity between brain areas, we obtain p-values for each identified connectivity epoch using the excursion test of ventura2005statistical, as follows. After computing p values $p^{(t,s)}$ for each precision entry (Eq. 15), we define a cluster as any contiguous set of entries whose p-values fall below $\frac{k_{\mathrm{BH}}}{n_{\mathrm{roi}}}\alpha_{\mathrm{BH}}$. No additional cluster forming threshold or minimum cluster size is imposed. For each cluster $k$ identified by the BH procedure, we calculate the test statistic:
@@
which is reminiscent of Fisher’s method for testing the global significance of multiple hypotheses (fisher1925statistical). Large values of $T_{k}$ provide evidence against cross-area connectivity in cluster $k$, so we calculate the corresponding p-value as $\int_{T_{k}}^{\infty}f_{T_{\max}}(t)~dt$, where $f_{T_{\max}}$ is the null distribution of $T_{\max}:=\max_{j}T_{j}$ under the global null hypothesis of no connectivity anywhere. We use $f_{T_{\max}}$ rather than the respective null distributions of each $T_{k}$ to control the family-wise type I error rather than the type I error for each cluster. We approximate $f_{T_{\max}}$ by the previous permutation bootstrap: for each permuted dataset $b=1,\ldots,B$, we estimate the cross-precision matrix and corresponding p-values, identify all clusters of p-values below $\frac{k_{\text{BH}}}{n_{\text{roi}}}\alpha_{\text{BH}}$, calculate the corresponding test statistics in Eq. 17, and let $T_{\max}^{b}$ be their maximum. The $B$ values $T_{\max}^{b}$ are samples from $f_{T_{\max}}$, which we use to approximate the p-value for cluster $k$ by the sampling proportion:
@@ 3.5 Locally Stationary State-space Model and Local Granger Causality
This state-space formulation is convenient to impose local stationarity on the latent time series, which we do by fitting stationary state-space models in moving windows of time. Local stationarity is justified because the functional connectivity within and between brain regions changes relatively slowly over time. The state-space formulation is also convenient to calculate the Granger causality between regions: $Z_{2}$ is said to Granger-cause $Z_{1}$ at time $t$ if some $\alpha_{12,s}^{(t)}$ are non-zero (ombao2022spectral) (and conversely if some $\alpha_{21,s}^{(t)}$ are non-zero). The coefficient of partial determination (partial $R^{2}$) between $\left(Z_{2}^{(t-d_{\mathrm{cross}})},\dots,Z_{2}^{(t-1)}\right)$ and $Z_{1}^{(t)}$, conditional on $Z_{1}^{(t-d_{\mathrm{auto}})},\dots,Z_{1}^{(t-1)}$, calculated as
@@
where $\tau_{2}$ is the maximum connection time from brain regions 2 to 1, $\tau_{1}\leq\tau_{2}\leq d_{\mathrm{cross}}$. We set $\tau_{2}=d_{\mathrm{cross}}$ by default, unless there is reason to consider shorter connection times. A plug-in estimator of $R^{2}_{2\rightarrow 1}(t)$ is easily obtained from the estimated covariance matrix of $\left(Z_{1}^{(t)},\dots,Z_{1}^{(t-d_{\mathrm{auto}})},Z_{2}^{(t-1)},\dots,Z_{2}^{(t-d_{\mathrm{cross}})}\right),$ without actually running the regressions (anderson1994model).
@@ 4 Results
We have introduced LaDynS to estimate the dynamic connectivity between two or more multivariate time series, and we proposed inference procedures to identify when connectivity is statistically significant. We apply LaDynS to experimental data in Section 4.4, but first we present the theoretical results on the equivalence between generative pCCA models and model-free CCA methods. In particular, Theorem 4.3 establishes an equivalence between the GENVAR version of multi-set CCA (kettenring1971canonical) and maximum likelihood applied to our dynamic pCCA. Next we examine its performance on simulated data that have properties similar to the experimental data. The data are simulated from the shared oscillatory driver model described in Section 4.2, which is unrelated to the LaDynS model. Section S3.2 contains results based on simulated data that are consistent with the LaDynS model. In Section 4.3, we compare the performance of LaDynS to other existing methods. The reproducible code scripts for the simulations and experimental data analyses are provided at github.com/HeejongBong/ladyns.
@@ 4.1 Equivalence between dynamic pCCA model and multiset CCA
We first revisit the connection between classical CCA and the probabilistic CCA formulation of bach2005probabilistic. Let $\{(X_{1[n]},X_{2[n]})\}_{n=1}^{N}$ be independent observations drawn from the joint distribution in Eq. 2, and denote the maximum likelihood estimates by $(\widehat{\beta}_{1},\widehat{\beta}_{2})$. The result below restates Theorem 2 of bach2005probabilistic.
Theorem 4.1 (bach2005probabilistic, Theorem 2).
@@
We now derive a similar connection between the multiset generalization of CCA introduced by kettenring1971canonical and the dynamic pCCA model in Eqs. 5 and 6. The proof is provided in Appendix S1.
@@ 4.2 LaDynS performance on simulated data from a shared oscillatory driver model with time delay
We used a probabilistic model of the inter-areal coherence in local field potentials to generate LFP time series with a dynamic lead-lag relationship between two brain areas. The shared oscillatory driver model with time delay in ombao2022spectral assumes that the coherence between two univariate LFP time series $L_{1}$ and $L_{2}$ is driven by a univariate latent oscillation $L_{0}$ at frequency $f_{0}$:
@@
where $\eta_{k}^{(t)}$ is regional baseline noise, and $\tau_{k}$ is the lead-lag from $L_{0}$ to $L_{k}$. This model generalizes the Synaptic-Source-Mixing (SSM) model of schneider2021mechanism, which was shown to be capable of modeling the dynamic cross-regional coherence observed in experimental LFP data between frontal and parietal cortices, as well as between LGN and the visual cortex.
For our simulation, we extended Eq. 26 to generate two sets of multi-dimensional LFP time series $L_{k}$ of dimension $d_{k}$, $k=1,2$, and we used three latent oscillations $L_{0,j}$, $j=1,2,3$, to allow non-stationary lead-lag relationships between the two brain areas. The resulting model was a special case of the Latent Dynamic Factor Analysis model in bong2020latent:
@@
where the factor loadings $\beta_{kj}\in\mathbb{R}^{d_{k}}$ were vectors of dimension $d_{k}=25$. We virtually arranged the $25$ components of $L_{1}^{(t)}$ and $L_{2}^{(t)}$ on $5\times 5$ regular two-dimensional arrays so we could create simulated data exhibiting spatial correlations similar to those in the experimental data in Section 4.4. To do that we let $\eta_{k}^{(t)}$ be spatially correlated Gaussian noise whose temporal power spectrum scales as $1/f^{\alpha}$ for $\alpha>1$. More specifically, for each temporal frequency $f$, the respective Fourier coefficients $\widehat{\eta}_{ki}(f)$ of region $k$ and electrode $i$ has spatial correlation: $\operatorname{Cov}[\widehat{\eta}_{ki}(f),\widehat{\eta}_{kj}(f)]=f^{-\alpha}\exp(-\frac{\mathrm{dist}^{2}(i,j)}{2\sigma_{\mathrm{spatial}}^{2}})$, where $\mathrm{dist}(i,j)$ was the distance between electrodes $i$ and $j$ on one array, $\alpha=1.4$, and $\sigma_{\text{spatial}}=0.8$. Next, we let the factor loadings $\beta_{kj}$ have components $\beta_{kji}=\gamma\cdot\exp(-\frac{\mathrm{dist}^{2}(i,p_{kj})}{2\sigma_{\mathrm{spatial}}^{2}})$, $i=1,\ldots d_{k}$, where $p_{kj}$ was a randomly chosen location on an array and $\gamma$ took values in $\{0.055,0.063,0.071,0.077,0.084,0.089\}$. This set spanned $\{0.5,0.66,0.84,1.0,1.16,1.34\}$ times the observed signal-to-noise ratio (SNR) in the experimental data in Section 4.4. More specifically, the observed data exhibited a signal to noise ratio (defined using signal power) of approximately 0.75. Finally, the three latent oscillations $L_{0,j}$ were set at $18$ Hz and designed to induce three epochs of lead-lag relationships between $L_{1}$ and $L_{2}$ around experimental times $80$, $200$ and $400$ ms; $L_{1}$ lead $L_{2}$ by $30$ ms during the first epoch, and $L_{1}$ lagged $L_{2}$ by $30$ ms during the other two, as depicted in Figure 5(a).
One simulated dataset consisted of $N=1000$ trials of $500$ ms long $d_{1}=25$ and $d_{2}=25$-dimensional time series $\{L_{1}^{(t)}\in\mathbb{R}^{d_{1}},L_{2}^{(t)}\in\mathbb{R}^{d_{2}}\}$ generated from Eq. 27 at sampling frequency $1000$ Hz. We filtered the simulated LFP recordings using the complex Morlet wavelet at frequency $f_{0}=18$ Hz and bandwidth $50$ ms, the same we applied to the data in Section 4.4, and collected the beta oscillation amplitude envelopes as the absolute values of the filtered signals. The complex Morlet wavelet is a complex sinusoidal with a Gaussian envelope (gabor1946theory), where the bandwidth refers to the Gaussian standard deviation. (The filtered amplitude at a given time $t$ is essentially the amplitude that would be obtained using a harmonic regression with cosine and sine damped by the Gaussian kernel centered at $t$; the wavelet formulation is computationally different and more efficient.) After downsampling the power envelopes to $100$ Hz, we applied LaDynS to the resulting data $X_{1}$ and $X_{2}$ with $T=50$ time points.
@@ 4.2.2 Results
Figures 5 and 6 used connection strength $\gamma=0.077$. We repeated the simulation for a range of $\gamma$ such that the SNR in the simulated data beta band ranged from zero to twice the estimated SNR of the experimental data in Section 4.4. Instead of showing graphs we summarized the performance of LaDynS using false cluster discovery and non-discovery rates (perone2004false) (FCDR, FCNR) in the estimated precision matrix. We defined a cluster of estimated nonzero precision elements to be falsely discovered if it contained no true effect, and a true cluster was deemed falsely non-discovered if no estimated cluster overlapped with it. We estimated these rates with the corresponding proportions in $60$ simulated datasets. They are plotted in Figure 7 against nominal significance levels of the excursion test between 0 and 10%, where the target FDR for the BH procedure was fixed at 5%. The FCDR remained below the nominal level across all tested significance values and connectivity intensities (all or almost all discovered clusters were true clusters) and for large enough $\gamma$, the FCNR was zero (all clusters were discovered).